Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi.


C = 5π; A = 20π

C = 10π; A = 20π

C = 10π; A = 25π

C = 5π; A = 25π


1
Expert's answer
2020-07-02T18:12:35-0400

d =10

"r = d\/2"

"С = 2\pi r =10\pi"

"A = \pi r^2 = 25\pi"

Answer: C = 10π; A = 25π